OpenStudy (diyadiya):

Two candidates attempts to solve a quadratic equation of form \(x^2+px+q=0\) One starts with a wrong value of p and find the roots to be 2 and 6 the other starts with a wrong value of q and finds the roots to be 2, -9 the correct roots are :

OpenStudy (anonymous):

Just put in the roots into the quadratic equations, you will get 4 equations in two variables, solve them to get the value of correct p and q. And then solve the quadratic equation to get the right roots.

OpenStudy (diyadiya):

Ok i'll try that

OpenStudy (anonymous):

By using the relation between roots and coefficient, -p=(2+6) p=-8 q=(2)(-9) = -18 x2-8x-18=0 solve this equation. (x+10)(x-2)=0 x=-10 or 2

OpenStudy (anonymous):

Oh Nice!

OpenStudy (diyadiya):

But they start with wrong value of p

OpenStudy (callisto):

I was thinking 'One starts with a wrong value of p and find the roots to be 2 and 6', that means q is correct, that is q =2x6 =12 ' the other starts with a wrong value of q and finds the roots to be 2, -9' that means p is correct -p = -9+2 p =7 just put the values into the equation and solve it.. Dunno if it is correct..

OpenStudy (diyadiya):

Yeah that's right :) @Callisto

OpenStudy (diyadiya):

-3 and -4 ?

OpenStudy (callisto):

That's what i've got.. :S

OpenStudy (diyadiya):

ok thats right :D Thanks everyone :)

OpenStudy (angela210793):

I think Vieta's formulas would work in here

OpenStudy (diyadiya):

What's Vieta's formula ?

OpenStudy (angela210793):

http://en.wikipedia.org/wiki/Vieta%27s_formulas